OpenStudy (anonymous):

PLZZ HELP PPLE square root of 2/3 is what? in fraction form

OpenStudy (anonymous):

.81649

OpenStudy (anonymous):

Is this is the fraction form @catamountz15

OpenStudy (anonymous):

lolz no

OpenStudy (anonymous):

It is already in fraction form so what more can you do here ??

OpenStudy (anonymous):

If you can, the sqrt(2/3) is just that. reducing it to a decimal answer is kind of rediculous but if you're required to do that and are just beginning in math i'd just use a calculator. There are several online ones. wolfram alpha being one of the best. just my two cents! Good luck!

OpenStudy (anonymous):

ehre are the choices

OpenStudy (anonymous):

Show them..

OpenStudy (anonymous):

@waterineyes sorry my bad did not see the in fraction form part. Il work it out in fraction form

OpenStudy (anonymous):

square root6 / 3 square root 2 / 3 2/3 1/ square roto 3

OpenStudy (anonymous):

No need to say sorry..

OpenStudy (anonymous):

yah no need to say sorry

OpenStudy (anonymous):

Is that second option same as the question @pitachips

OpenStudy (anonymous):

no its square root 2/ normal 3

OpenStudy (anonymous):

Is this the question ??

OpenStudy (anonymous):

Or second option ??

OpenStudy (anonymous):

\[\sqrt{\frac{2}{3}}\] Or: \[\frac{\sqrt{2}}{3}\] ?? What is the question actually ??

OpenStudy (anonymous):

the first one is the question the second one is the second answer choice HOW DO YOOU DO THAT????????

OpenStudy (anonymous):

Multiply 3 to both numerator and denominator: \[\sqrt{\frac{2}{3}} \implies \sqrt{\frac{2 \times 3}{3 \times 3}} \implies \frac{\sqrt{2 \times 3}}{\sqrt{3 \times 3}} = \frac{\sqrt{6}}{3}\]

OpenStudy (anonymous):

question is simplify this \[\sqrt{2/3}\] anwer choices are........\[\sqrt{6}\]/3 \[\sqrt{2}\]/3 2/3 1/\[\sqrt{3}\]

OpenStudy (anonymous):

In equation editor use : frac{}{} note the brackets.. suppose you want to write 2/3 then in equation editor write frac{2}{3}

OpenStudy (anonymous):

For writing sqrt{6}{3} try this in your equation editor: frac{sqrt{6}}{3}

OpenStudy (anonymous):

woooooow thnx leet me fix the answer choices then

OpenStudy (anonymous):

can you wait one min

OpenStudy (anonymous):

\[\huge \boxed {\color{green}{\textbf{Welcome To Openstudy..}}}\]

OpenStudy (anonymous):

\[\color{green}{\sqrt{\frac{2}{3}}} \implies \color{blue}{\sqrt{\frac{2 \times 3}{3 \times 3}}} \color{red}{\implies \frac{\sqrt{2 \times 3}}{\sqrt{3 \times 3}}} \implies \color{orange}{\frac{\sqrt{6}}{3}}\]

OpenStudy (anonymous):

\[\frac{\sqrt{6}}{3} \] \[\frac{\sqrt{2}}{3}\] 2/3 \[\frac{1}{\sqrt{5}}\]

OpenStudy (anonymous):

Well done.. you can edit 2/3 as well frac{2}{3}

OpenStudy (anonymous):

k BUT WHATS THE ANSWER GOT CARRIED AWAY im going to fail my test

OpenStudy (anonymous):

What I have written in orange above ??

OpenStudy (anonymous):

OH THNXX!!! nooooooo i got a 70 i got the last 3 wrong there are only 10 questions

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

haha thanks